# Value-Incentivized Preference Optimization: A Unified Approach to Online and Offline RLHF

### v3

@@ 1.2 Related work
Since the introduction of the original RLHF framework, there have been many proposed simplifications of the preference alignment procedure and attempts to improve performance, including but not limited to SLiC (Zhao et al.,, 2023), GSHF (Xiong et al.,, 2023), DPO (Rafailov et al.,, 2023), and its variants, such as Nash-MD (Munos et al.,, 2023), IPO (Azar et al.,, 2024), OAIF (Guo et al.,, 2024), SPO (Swamy et al.,, 2024), SPIN (Chen et al.,, 2024), GPO (Tang et al.,, 2024), and DPOP (Pal et al.,, 2024). These methods can roughly be grouped into online and offline variants, depending on whether preference data is collected before training (offline) or by using the current policy during training (online).
@@
The principles of optimism and pessimism are typically implemented via constructing confidence intervals or posterior sampling, which have been demonstrated to be provably efficient in tabular settings (Jin et al.,, 2018; Shi et al.,, 2022). Yet, these approaches have had limited success in conjunction with deep learning architectures (Gawlikowski et al.,, 2021), and many empirical heuristics in turn lack theoretical validation (Kumar et al.,, 2020). Notwithstanding, alternative regularization schemes are developed for general function approximation settings with theoretical guarantees, such as Bellman-consistent pessimism for offline RL (Xie et al.,, 2021), and reward-biased exploration for online RL (Mete et al.,, 2021; Liu et al., 2024a, ). VPO draws inspiration from reward-biased exploration (Kumar and Becker,, 1982; Liu et al.,, 2020; Liu et al., 2024a, ; Hung et al.,, 2021; Mete et al.,, 2021) in the standard online RL literature, but significantly broadens its scope to the offline setting and RLHF for the first time.
@@
Since posting the initial version of this paper on arXiv, we discovered several concurrent work that also appeared online around the same time proposing similar regularization techniques as ours to encourage optimism (resp. pessimism) for online (resp. offline) RLHF (Zhang et al.,, 2024; Xie et al.,, 2024; Liu et al., 2024b, ). Despite slightly different forms, the algorithms studied in Zhang et al., (2024); Xie et al., (2024); Liu et al., 2024b can be interpreted as adopting different choices of the calibration policy in VPO. In the context of online RLHF, Zhang et al., (2024) empirically studied a similar algorithm as the proposed online VPO under the contextual bandit formulation of RLHF; Xie et al., (2024) provided finite-time regret analysis of a similar algorithm for the token-level MDP formulation with general function approximation, which extends to general deterministic contextual MDPs as well. In the context of offline RLHF, Liu et al., 2024b studied a similar algorithm as the proposed offline VPO and provided a sample complexity analysis under the contextual bandit formulation, yet focusing on general function approximation and different assumptions.
@@ 5 Experiments
In this section, we evaluate the proposed VPO on both synthetic multi-armed bandit (MAB), and RLHF for LLMs, in online and offline settings.
5.1 Synthetic multi-armed bandits
We evaluate the proposed methods on a synthetic dataset of size $|\mathcal{X}|=1$ and $|\mathcal{Y}|=10$. We set $\pi_{\text{ref}}=\pi_{\mathsf{b}}=\mathcal{\pi}_{\text{cal}}$, where $\pi_{\text{ref}}=\mathsf{softmax}(\theta_{\text{ref}})$ with $\theta_{\text{ref}}(x,y)$ sampled i.i.d. from $\mathcal{N}(0,1)$. The ground truth reward $r^{\star}$ is randomly generated i.i.d. according to $r^{\star}(x,y)\sim U([0,1])$. We approximately solve the optimization problems by performing $20$ AdamW optimization steps with learning rate $0.01$ and weight decay rate $0.01$ in every iteration for the online setting and $1000$ steps for the offline setting.
We plot the average results over 10 independent runs in Figure 1. As demonstrated in the left panel of Figure 1, an appropriate choice of $\alpha$ allows our method to outperform the model-based MAB with MLE baseline in the long-term performance of cumulative regret, at the cost of slightly increased cumulative regret in the first 100 iterations. This highlights the effectiveness of the VPO in achieving more principled exploration-exploitation trade-off. For the offline setting, the right panel of Figure 1 demonstrates that the performance of both MLE-MAB and VPO improves as the number of offline data increases. However, VPO achieves a consistently lower sub-optimality gap compared with that of MLE-MAB.
@@ 5.1 Synthetic multi-armed bandits
Figure 1: The cumulative regret v.s. number of iterations plot (left panel) and sub-optimality gap v.s. number of data plot (right panel) of VPO and MLE-MAB methods in the online and offline settings, respectively.
@@ 5.2 RLHF for LLMs
In this setting, we evaluate the performance of VPO on the TL;DR task (Stiennon et al.,, 2020). We prepare the prompts dataset by extracting the input prompts from the preference data. Recall we are evaluating the algorithm performance in online setting, we only compare to the online RLHF baselines (Guo et al.,, 2024) for fairness. We adopt PaLM2  (Anil et al.,, 2023) as the language model and also the LLM annotator. We conduct VPO and online DPO to the same PaLM2-XXS as the policy, which is initialized by supervised finetuning, denoted as SFT model. We exploit another PaLM2-XS model as the LLM annotator to provide online feedbacks. Similar to  (Guo et al.,, 2024), we use Detailed 0-shot prompt from Lee et al., (2023). The prompts we used and how we get preference scores are detailed in Appendix C. We emphasize our algorithm is agnostic to human or AI feedback.
As a sanity check, we track the win rate of VPO and online DPO against the SFT baseline on TL;DR during training in Figure 5.2. For ablation purpose, we varies the exploration weight $\alpha=\{0.01,0.1\}$ in the optimistic VPO. One significant observation is that although all the online RLHF algorithms follow the increase trend, the win-rate against SFT of the optimistic VPO has larger oscillation, comparing to online DPO. And the oscillation reduces, with $\alpha$ diminishing. Our conjecture is that this behavior is encouraged by the optimistic term in VPO, for collecting more unexplored data, which may delay the learning due to the diversity in data. However, as the learning proceeds, the proposed VPO outperforms the competitors, because of the coverage of the collected data.
@@
To demonstrate the advantages of optimistic VPO in online setting more directly, we evaluate the win/tie/loss rate against online DPO head-to-head, as shown in Figure 5.2. This clearly shows that the optimistic VPO achieves better performances with larger exploration preference, and thus, consolidates our conclusion that i), the simple value-incentivized term makes the exploration practical without uncerntainty estimation; and ii), exploration is potentially beneficial for better model.
@@ Acknowledgement
The work of S. Cen and Y. Chi is supported in part by the grants ONR N00014-19-1-2404, NSF DMS-2134080, CCF-2106778 and CNS-2148212. S. Cen is also gratefully supported by Wei Shen and Xuehong Zhang Presidential Fellowship and JP Morgan AI Research Fellowship.
@@ A.1 Proof of Theorem 1
$=\sum_{t=1}^{T}\big{\langle}{W(r^{(t)}),X(r^{(t)})}\big{\rangle}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\}\leq 1\}$
$\quad\quad\quad+\sum_{t=1}^{T}\big{\langle}{W(r^{(t)}),X(r^{(t)})}\big{\rangle}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\}>1\},$ (45)
@@
    $\sum_{t=1}^{T}\big{\langle}{W(r^{(t)}),X(r^{(t)})}\big{\rangle}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\}\leq 1\}$
    $\leq\sum_{t=1}^{T}\|W(r^{(t)})\|_{\Sigma_{t}}\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\}\leq 1\}$
@@
    $\sum_{t=1}^{T}\big{\langle}{W(r^{(t)}),X(r^{(t)})}\big{\rangle}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\}>1\}$ $\leq C\sum_{t=1}^{T}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\}>1\}$
@@ Appendix C Experimental details
C.1 Offline setting
For the offline setting experiments, we adopt instruction tuned models, Llama2-13b-chat and Flan-T5-xl as the base models. To prompt these models, we prepend the question with
@@ C.1 Offline setting
The question is structured in a way that the multiple choices are shown as alphabets (letters) within parenthesis.
We set $\mathcal{\pi}_{\text{cal}}$ to the empirical distribution of the ground truth answer which is known to us. Based on preliminary experiments, we set $\beta$ as 0.1 in DPO and $\tau$ as 1.0 in IPO. For VPO, we experiment with moving $\alpha$ from 0.01 to 10, choosing 1 for the reported results for Flan-T5-xl results. For experiments on Llama2-13b-chat, we also set $\alpha$ to 1.
For both models, we train the base models with different algorithms DPO, VPO and IPO for 3000 steps and report the accuracy of the performance on the ARC-challenge test data set after every 500 steps. The training for Llama2-13b-chat model on 128 TPU-v4 takes around 2hrs and for Flan-T5-xl on 64 TPU-v3 takes 1 hour.
C.2 Online setting

### v4

@@ 1.2 Related work
Since the introduction of the original RLHF framework, there have been many proposed simplifications of the preference alignment procedure and attempts to improve performance, including but not limited to SLiC (Zhao et al.,, 2023), GSHF (Xiong et al.,, 2023), DPO (Rafailov et al.,, 2023), and its variants, such as Nash-MD (Munos et al.,, 2023), IPO (Azar et al.,, 2024), OAIF (Guo et al.,, 2024), SPO (Swamy et al.,, 2024), SPIN (Chen et al.,, 2024), WIND (Yang et al.,, 2024), GPO (Tang et al.,, 2024), and DPOP (Pal et al.,, 2024). These methods can roughly be grouped into online and offline variants, depending on whether preference data is collected before training (offline) or by using the current policy during training (online).
@@
The principles of optimism and pessimism are typically implemented via constructing confidence intervals or posterior sampling, which have been demonstrated to be provably efficient in tabular settings (Jin et al.,, 2018; Shi et al.,, 2022). Yet, these approaches have had limited success in conjunction with deep learning architectures (Gawlikowski et al.,, 2021), and many empirical heuristics in turn lack theoretical validation (Kumar et al.,, 2020). Notwithstanding, alternative regularization schemes are developed for general function approximation settings with theoretical guarantees, such as Bellman-consistent pessimism for offline RL (Xie et al.,, 2021), and reward-biased exploration for online RL (Mete et al.,, 2021; Liu et al., 2024a, ) and Markov games (Liu et al., 2024a, ; Yang et al.,, 2025). VPO draws inspiration from reward-biased exploration (Kumar and Becker,, 1982; Liu et al.,, 2020; Liu et al., 2024a, ; Hung et al.,, 2021; Mete et al.,, 2021) in the standard online RL literature, but significantly broadens its scope to the offline setting and RLHF for the first time.
@@
Since posting the initial version of this paper on arXiv, we discovered several concurrent work that also appeared online around the same time proposing similar regularization techniques as ours to encourage optimism (resp. pessimism) for online (resp. offline) RLHF (Zhang et al., 2024a, ; Xie et al.,, 2024; Liu et al., 2024b, ). Despite slightly different forms, the algorithms studied in Zhang et al., 2024a ; Xie et al., (2024); Liu et al., 2024b can be interpreted as adopting different choices of the calibration policy in VPO. In the context of online RLHF, Zhang et al., 2024a empirically studied a similar algorithm as the proposed online VPO under the contextual bandit formulation of RLHF; Xie et al., (2024) provided finite-time regret analysis of a similar algorithm for the token-level MDP formulation with general function approximation, which extends to general deterministic contextual MDPs as well. In the context of offline RLHF, Liu et al., 2024b studied a similar algorithm as the proposed offline VPO and provided a sample complexity analysis under the contextual bandit formulation, yet focusing on general function approximation and different assumptions.
@@ 5 Experiments
In this section, we evaluate the proposed VPO on both synthetic bandits and RLHF for LLMs, in online and offline settings.
5.1 Synthetic bandit problems
We evaluate the proposed methods on two synthetic problems: i) a multi-armed bandit problem, and ii) a linear contextual bandit problem.
Multi-armed bandit (MAB) problem.
In this scenario, we set $|\mathcal{X}|=1$ and $|\mathcal{Y}|=10$. For each $y\in\mathcal{Y}$, the ground truth reward $r^{\star}(x,y)$ is randomly generated i.i.d. from a uniform distribution $U([0,1])$. The policy is parameterized by $\pi_{\theta}(\cdot|x)=\mathsf{softmax}(\theta(x,\cdot))$, where $\theta\in\mathbb{R}^{10}$. The reference policy $\pi_{\text{ref}}$ is set to $\pi_{\theta_{\text{ref}}}$ with $\theta_{\text{ref}}(x,y)$ sampled i.i.d. from $U([0,1])$.
Linear contextual bandit problem.
Here, we set $\mathcal{X}=\mathbb{R}^{2}$, $|\mathcal{Y}|=50$. For each $(x,y)$ pair, the ground truth reward is given by $r^{\star}(x,y)=\left\langle\phi(x,y),\theta^{\star}\right\rangle$, where $\theta^{\star}\in\mathbb{R}^{10}$ is randomly sampled from $U([0,1])$, and the feature vector $\phi(x,y)$ is the output of the hidden layer of a fixed two-layer MLP, with the input given by the concatenation of $x$ and the one-hot encoding of $y$. The activation function is set to tanh. The context vector $x$ is drawn from standard normal distribution. We focus on log-linear policy class $\pi_{\theta}(\cdot|x)=\mathsf{softmax}(\langle\theta,\phi(x,\cdot)\rangle)$, and set $\pi_{\text{ref}}=\pi_{\theta_{\text{ref}}}$ with $\theta_{\text{ref}}(x,y)$ sampled i.i.d. from $U([0,1])$.
For both problem we set $\pi_{\text{ref}}=\pi_{\mathsf{b}}=\mathcal{\pi}_{\text{cal}}$ and use mini-batch sample of size 5 in every iteration. We approximately solve the optimization problems by performing $20$ AdamW optimization steps with learning rate $0.01$ and weight decay rate $0.01$ in every iteration for the online setting and $1000$ steps for the offline setting. We set $\beta=5$ for the online linear contextual bandit problem and $\beta=1$ for all other experiments to better illustrate the performance differences.
We plot the average results over 50 independent runs for both experiments in Figure 1 and Figure 2. As demonstrated in Figure 1, an appropriate choice of $\alpha$ allows our method to outperform the model-based MAB with MLE baseline in the long-term performance of cumulative regret, at the cost of slightly increased cumulative regret in the first 100 iterations. This highlights the effectiveness of the VPO in achieving more principled exploration-exploitation trade-off. For the offline setting, Figure 2 demonstrates that the performance of both MLE-MAB and VPO improves as the number of offline data increases. However, VPO achieves a consistently lower sub-optimality gap compared with that of MLE-MAB.
@@ 5.1 Synthetic bandit problems
Figure 1: The cumulative regret v.s. number of iterations plot of VPO and MLE-MAB methods in the online MAB problem (left panel) and online linear contextual bandit problem (right panel), respectively.
Refer to caption Refer to caption
Figure 2: The sub-optimality gap v.s. number of data plot of VPO and MLE-MAB methods in the offline MAB problem (left panel) and offline linear contextual bandit problem (right panel), respectively.
@@ 5.2 RLHF for LLMs
For online setting, we conduct two distinct experimental setups. The first, referred to as Buffer, adopts the experimental setup in Online AI Feedback (OAIF)  (Guo et al.,, 2024), with an additional buffer used to sample the data for the exploration part of the VPO loss. The second setup, referred to as Iterative, adopts the experimental setup in  (Zhang et al., 2024a, ), relying on an online iterative training framework.
Buffer.
In these set of experiments, we adopt OAIF experimental setup  (Guo et al.,, 2024) where the preference data is gathered by online sampling from the policy and labeled through online feedback. We also introduce a buffer that stores the labeled preferences and is used to sample the data for the exploration term in the VPO loss. We adopt PaLM2-XXS language model  (Anil et al.,, 2023) as policy, initialized by supervised finetuning, denoted as SFT model. We exploit another PaLM2-XS model as the LLM annotator to provide online feedback. We evaluate the performance of optimistic VPO and compare its performance to Online DPO  (Guo et al.,, 2024). We choose TL;DR task (Stiennon et al.,, 2020) and extract its prompts for the input of preference data. Similar to  (Guo et al.,, 2024), we use Detailed 0-shot prompt from Lee et al., (2023). The prompts we used and how we get preference scores are detailed in Appendix C. We emphasize our algorithm is agnostic to human or AI feedback.
As a sanity check, we track the win rate of VPO and Online DPO against the SFT baseline on TL;DR during training in Figure 5.2. For ablation purpose, we vary the exploration weight $\alpha=\{0.01,0.1\}$ in the optimistic VPO. One significant observation is that although all the online RLHF algorithms follow the increase trend, the win-rate against SFT of the optimistic VPO has larger oscillation, comparing to Online DPO. And the oscillation reduces, with $\alpha$ diminishing. Our conjecture is that this behavior is encouraged by the optimistic term in VPO, for collecting more unexplored data, which may delay the learning due to the diversity in data. However, as the learning proceeds, the proposed VPO outperforms the competitors, because of the coverage of the collected data.
@@
To demonstrate the advantages of optimistic VPO in online setting more directly, we evaluate the win/tie/loss rate against Online DPO head-to-head, as shown in Figure 5.2. This clearly shows that the optimistic VPO achieves better performances with larger exploration preference, and thus, consolidates our conclusion that i), the simple value-incentivized term makes the exploration practical without uncertainty estimation; and ii), exploration is potentially beneficial for better model.
Iterative.
We further evaluate the performance of VPO on standard benchmarks AlpacaEval 2.0 (Dubois et al., 2024b, ; Dubois et al., 2024a, ) and MT-Bench (Zheng et al.,, 2024), using similar experimental setup in recent literature (Zhang et al., 2024b, ; Wu et al.,, 2024). We use UltraFeedback [^1] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized (Cui et al.,, 2023) as our training dataset which contains around 61k preference pairs of single-turn conversations. We split the 61k prompts into four chunks and follow an iterative training approach. We choose Zephyr-7B-SFT [^2] https://huggingface.co/HuggingFaceH4/mistral-7b-sft-beta as our LLM model. We follow the best hyperparameters setup found in Zhang et al., 2024b and compare the results of VPO to DPO reported therein. We first conduct a single iteration of standard DPO training on the first portion of the training data, referred to as Zephyr-7B-DPO in  1. We then perform 3 iterations of VPO, each iteration on a different data portion, while using online AI feedback provided by PairRM (Jiang et al.,, 2023) in between. Further details of our experiments are explained in Appendix C.
The evaluation results are summarized in Table 1, filling out the baselines based on Zhang et al., 2024b . As could been seen, VPO significantly improves the performance of the base model Zephyr-7B-SFT by 14.52, achieving the highest length-controlled (LC) Win Rate on the AlpacaEval 2.0 benchmark, beating DPO. This result, 22.53, is competitive to much larger models such as much Yi-34B-Chat, 27.19, and Llama-3-70B-Instruct, 33.17. Additionally, VPO shows significant improvement on MT-Bench compared to the base model Zephyr-7B-SFT with the increase of 2.32 while beating DPO. We further report the results of other iterative post-training algorithms, such as SPIN (Chen et al.,, 2024), DNO (Rosset et al.,, 2024), and SPPO (Wu et al.,, 2024) and show that even though VPO is trained on a weak base model, it achieves comparable results to these baselines. Granular views on a radar chart can be found in Appendix C.
Model AlpacaEval 2.0 MT-Bench
LC Win Rate Win Rate Avg. Len Avg 1st Turn 2nd Turn
Zephyr-7B-SFT 8.01 4.63 916 5.30 5.63 4.97
Zephyr-7B-DPO 15.41 14.44 1752 7.31 7.55 7.07
DPO Iter 1 (Zephyr) 20.53 16.69 1598 7.53 7.81 7.25
DPO Iter 2 (Zephyr) 22.12 19.82 1717 7.55 7.85 7.24
DPO Iter 3 (Zephyr) 22.19 19.88 1717 7.46 7.85 7.06
VPO Iter 1 (Zephyr) 22.53 19.09 1638 7.50 7.76 7.24
VPO Iter 2 (Zephyr) 21.84 18.78 1663 7.62 7.93 7.32
VPO Iter 3 (Zephyr) 22.15 19.58 1713 7.61 8.01 7.21
SPIN 7.23 6.54 1426 6.54 6.94 6.14
Orca-2.5-SFT 10.76 6.99 1174 6.88 7.72 6.02
DNO (Orca-2.5-SFT) 22.59 24.97 2228 7.48 7.62 7.35
Mistral-7B-Instruct-v0.2 19.39 15.75 1565 7.51 7.78 7.25
SPPO (Mistral-it) 28.53 31.02 2163 7.59 7.84 7.34
Yi-34B-Chat 27.19 21.23 2123 7.90 - -
Llama-3-70B-Instruct 33.17 33.18 1919 9.01 9.21 8.80
GPT-4 Turbo (04/09) 55.02 46.12 1802 9.19 9.38 9.00
Table 1: Results on AlpacaEval 2.0 and MT-Bench.
@@ Acknowledgement
The work of S. Cen, T. Yang and Y. Chi is supported in part by the grants ONR N00014-19-1-2404, NSF DMS-2134080, CCF-2106778 and CNS-2148212. S. Cen is also gratefully supported by Wei Shen and Xuehong Zhang Presidential Fellowship and JP Morgan AI Research Fellowship.
@@ A.1 Proof of Theorem 1
$=\sum_{t=1}^{T}\big{\langle}{W(r^{(t)}),X(r^{(t)})}\big{\rangle}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\leq 1\}$
$\quad\quad\quad+\sum_{t=1}^{T}\big{\langle}{W(r^{(t)}),X(r^{(t)})}\big{\rangle}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}>1\},$ (45)
@@
    $\sum_{t=1}^{T}\big{\langle}{W(r^{(t)}),X(r^{(t)})}\big{\rangle}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\leq 1\}$
    $\leq\sum_{t=1}^{T}\|W(r^{(t)})\|_{\Sigma_{t}}\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}\leq 1\}$
@@
    $\sum_{t=1}^{T}\big{\langle}{W(r^{(t)}),X(r^{(t)})}\big{\rangle}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}>1\}$ $\leq C\sum_{t=1}^{T}\mathbf{1}\{\|X(r^{(t)})\|_{\Sigma_{t}^{-1}}>1\}$
@@ Appendix C Experimental details
C.1 RLHF for LLMs — offline setting
For the offline setting experiments, we adopt instruction tuned models, Llama2-7b-chat, Llama2-13b-chat and Flan-T5-xl as the base models. To prompt these models, we prepend the questions in ARC-Challenge task (Clark et al.,, 2018) with
@@ C.1 RLHF for LLMs — offline setting
The question is structured in a way that the multiple choices are appended with alphabets (letters) within parenthesis to the question. As an example:
Question: George wants to warm his hands quickly by rubbing them. Which skin surface will produce the most heat? Choices: (A)dry palms (B)wet palms (C)palms covered with oil (D)palms covered with lotion
We set $\mathcal{\pi}_{\text{cal}}$ to the empirical distribution of the ground truth answer which is known to us. Based on preliminary experiments, we set $\beta$ as 0.1 in DPO and $\tau$ as 1.0 in IPO. For VPO, we experiment with moving $\alpha$ from 0.01 to 10, choosing 1 for the reported results. For all models, we train the base models with different algorithms DPO, VPO and IPO for 3000 steps and report the accuracy of the performance on the ARC-challenge test data set after every 500 steps. The training for Llama2-13b-chat model on 128 TPU-v4 takes around 2hrs and for Flan-T5-xl on 64 TPU-v3 takes 1 hour.
C.2 RLHF for LLMs — online setting
Buffer.
@@ C.2 RLHF for LLMs — online setting
Iterative.
The UltraFeedback data (Cui et al.,, 2023) contains around 61k preference pairs of single-turn conversations. We divide this data set into 4 chunks. We use the first chunk to train a DPO model which we refer to as Zephyr-7B-DPO in Table 1. We use the remaining 3 chunks to train consecutive iterations of DPO and VPO, using the checkpoint from the previous iteration to initialize the policy for the current iteration. For each iteration, the prompts from the data are extracted and a new answer is sampled from the policy. We label the data using online AI feedback provided by PairRM (Jiang et al.,, 2023), using the similar ranking procedure as (Zhang et al., 2024b, ) where the new sample is ranked against $y_{w}$ and $y_{l}$ from the data. For VPO, we also adopt another ranking process where we sample two answer from the policy which are ranked against each other. We also set $\mathcal{\pi}_{\text{cal}}$ to the empirical distribution of the negative answer pairs $(x,y_{-})$ collected by the policy. We report the results of the best checkpoint of these two ranking procedures in Table 1. All experiments are conducted on 16xA100 GPUs.
Figure 5: Radar chart of MT-Bench results for Zephyr-7B.